# ShapeGraFormer: GraFormer-Based Network for Hand-Object Reconstruction from a Single Depth Map

### v1

@@ 3.3 VoxelNet: 3D Voxelized Shape
3.4 ShapeGraFormer
To obtain topologically-coherent hand-object mesh representation, from the voxelized hand and object predictions obtained from $VoxelNet$, we train a new $ShapeGraFormer$ Graformer2022 . The $ShapeGraFormer$ includes Graph Convolutional layers and Multi-headed Attention layers in order to convert depth features into a pose and a shape for both the hand and the object. Although it was originally designed to lift 2D poses to 3D, GraFormer’s design which combines the advantages of Graph Convolutional Networks and Transformers making it capable of effectively solving any problem that can be represented as a graph. Both the hand and the spherical objects have consistent topology and hence can be represented as a graph. Therefore, in our method, we utilize three separate GraFormers, namely: hand GraFormer, object GraFormer, and refinement GraFormer, with two feature extractors: a hand feature extractor and an object feature extractor. as described in the following subsection.
@@ 3.4 ShapeGraFormer
3.4.1 Feature Extractor and Graph Initialization
For the graph initialization, we use the outputs of the $VoxelNet$ and the raw voxelized depth as shown in Figure 1. Namely, we extract featuremaps $\mathcal{F}_{V}$ from the $VoxelNet$ and pass them through an MLP that converts them into a feature vector of size $256$. Additionally, we utilize a simple 3D CNN to extract features from the voxelized shapes and reduce them to $128$ features. Furthermore, a 3D Max Pool layer reduces the size of the raw voxelized depth from $44\times 44\times 4$ 44 to $11\times 11\times 1$ 11 creating a $1331$ feature vector. These features are then combined to create a hand feature vector $\mathcal{F}^{H}_{1715}$ of size $1715$ to initialize the graph vertices as shown in Figure 1. The same operation is repeated for the object’s vertices resulting in $\mathcal{F}^{O}_{1715}$.
3.4.2 Positional Embedding
In order to generate distinct features for each vertex in the graph, we propose a positional embedding layer that converts the vertices of the template meshes for both the hand and object into positional vectors $\mathcal{E}_{i}^{p}$ of the same size as the feature vector where i is the index of the ith vertex in the combined hand-object graph. The positional vectors $\mathcal{E}_{i}^{p}$ are then accumulated with the shared feature vectors $\mathcal{F}^{H}_{1715}$ and $\mathcal{F}^{O}_{1715}$ depending on whether the ith vertex belongs to the hand or the object in order to create a unique representation for each vertex. For the hand mesh, we adopt the default MANO hand as the template, while for the object mesh, we use the sphere that is utilized for deforming objects, see Section 3.1.
3.4.3 GraFormer Details
Each GraFormer consists of five consecutive GraAttention components followed by Chebyshev Graph Convolutional layers as described in Graformer2022 and shown in Figure 3. A GraAttention component consists of a Multi-headed Attention layer where the hidden dimension is $128$ and the number of heads is $4$, following the ablation study mentioned in Graformer2022 where they studied the impact of hidden dimension size and number of layers in the GraFormer on pose estimation. Compared to normal Transformers, the last fully-connected layer of the Multi-headed Attention in the GraFormer is a graph convolutional layer, not a feed-forward layer. The GraFormer also contains an input layer that maps the feature vector to the hidden dimension size and an output layer that maps the hidden dimension into the corresponding 3D coordinate value for each vertex in the graph. To initialize the adjacency matrix of the graph layers, we use the mesh faces of the MANO model along with the faces of the spherical mesh.
@@ 3.4.3 GraFormer Details
3.4.4 Refinement GraFormer
To enhance the realistic appearance and correct minor shape-related artifacts in the predicted object shapes, we add an additional GraFormer for refinement. We utilize the initial shape produced by the Hand and Object GraFormer as input to another positional embedding layer while we use $\mathcal{F}^{H}_{1715}$ and $\mathcal{F}^{O}_{1715}$ to initialize the new graph. In the case of using a refinement GraFormer during training, we get $2$ separate shapes from the $ShapeGraFormer$ and apply the same loss function mentioned in 3.4 on both of them.
3.5 Training and Implementation Details
We train all the components of our network on fully annotated public hand-object pose and shape datasets, namely, HO-3D hampali2020honnotate and DexYCB dexycb . For training, we use the Adam adam optimizer with a learning rate set to .001.001.001.001. For improved convergence, we train $PoseNet$ separately, then fix the obtained weights to train the remaining network components. All learning and inference are implemented in PyTorch and are conducted on an NVIDIA A100 GPU.
4 Experiments
In this section, we quantitatively and qualitatively evaluate the effectiveness of our 3D hand-object reconstruction pipeline on two popular datasets, namely HO-3D dataset (version v2 and v3) hampali2020honnotate and DexYCB dexycb . We additionally compare our pipeline with state-of-the-art approaches in Section 4.2. For quantitative evaluation and comparisons, we use the following two metrics: (i) the average 3D joint location error and (ii) the mean vertex location error over all test frames.
4.1 Datasets
@@
The HO-3D dataset hampali2020honnotate , is a publicly available dataset with 3D pose annotations for hands interacting with objects captured from third-person views. The dataset has multiple versions and we report the results on v2 and v3 of the dataset. For HO-3D (v3), the training set $\mathcal{D}:=\{\mathcal{D}^{train},\mathcal{D}^{valid}\}$ contains annotations for $|\mathcal{D}^{train}|=71,662$ and $|\mathcal{D}^{valid}|=10,927$ images and a total of $55$ sequences and $10$ different objects from the YCB-Video dataset Xiang2018 , $9$ for the training set and $1$ (unseen) for the evaluation set. The evaluation set $\mathcal{D}^{eval}$ comprises $13$ sequences with a total of $|\mathcal{D}^{eval}|=20,137$ frames addressing challenging scenarios: namely, (i) $3$ sequences with ($2$) seen objects and seen hands, (ii) $5$ sequences with ($1$) seen object but unseen hands, and (iii) $5$ sequences with seen hands but $1$ unseen object. Hands in the evaluation set are just annotated with the wrist coordinates, while the full hand is not annotated. Object pose and shape are annotated over all available sets.
@@
To extend our evaluation, we also train our network on the DexYCB dataset dexycb . DexYCB contains hand pose and shape and 6D object pose annotations for  $582k$ frames recorded on $10$ different subjects, using $20$ different objects and $8$ views. We use the S1 evaluation setup as specified by the authors where $\mathcal{D}^{valid}$ contains $1$ unseen subject and $\mathcal{D}^{eval}$ contains $2$ unseen subjects and $\mathcal{D}^{train}$ contains 7 subjects and all different objects. The exact split sizes are: $|\mathcal{D}^{train}|=407,088$, $|\mathcal{D}^{valid}|=58,592$, and $|\mathcal{D}^{eval}|=116,288$.
@@
4.2 Evaluation
In this section, we evaluate hand-object shape and pose reconstruction and compare it to state-of-the-art approaches in challenging scenarios.
@@ 4.2 Evaluation
We compare our work on HO-3D quantitatively with six different methods. The work of Hasson et al. hasson2019learning and THOR-Net  Aboukhadra_2023_WACV are the most related to ours in terms of the goals. However, as they focus on RGB inputs, comparisons are made up to scale. Malik et al. HandVoxNet2020 on the other hand, is based on depth inputs and has a comparable voxel-based network pipeline to our design. However, they ignore the presence of an object and reconstruct hands in isolation. We include the representative RGB-based hand reconstruction approach of Hampali et al. hampali2020honnotate , and ArtiBoost  yang2022artiboost for completeness. For a fair comparison, all methods have been (re-)trained on the HO-3D dataset. The first two methods provide publicly available results (hand-only), which we report in Table 1. We re-implemented HandVoxNet following the authors’ instructions and trained all the network components on HO-3D. In addition to that, we also report the root-relative pose estimation error in Table 2 and compare it to two benchmark methods mentioned by the DexYCB authors. We also show qualitative samples for hand reconstruction from DexYCB in Figure 6.
@@
Quantitative evaluation on hand pose and shape reconstruction, shows that our method outperforms existing methods, see Table 1 and Figures 4, 5.
Method Joint err. (cm) Mesh err. (cm)
Hampali et al. hampali2020honnotate 8.42 8.34
Hasson et al. hasson2019learning 4.97 4.96
HandVoxNet HandVoxNet2020 2.77 2.67
HandVoxNet++ HandVoxNet++2021 2.46 2.70
THOR-Net Aboukhadra_2023_WACV 2.63 2.63
@@
ArtiBoost yang2022artiboost 2.26 10.95
THOR-Net Aboukhadra_2023_WACV 2.56 2.37
@@
Table 1: Quantitative results on the evaluation set $\mathcal{D}^{eval}$ of HO-3D hampali2020honnotate dataset. Upper table: HO-3D (v2), lower table: HO-3D (v3).
We achieve a minimum pose and shape prediction improvement respectively of $0.27$ cm and $0.43$ cm over the state-of-the-art. The experiments show that training the network for hand reconstruction alone without object reconstruction improves hand reconstruction and that the refinement stage has no impact on hand reconstruction as shown in Table 1. However, joint training for the hand and the object outperforms other methods. Figure 5 shows our results on a few frames of the challenging evaluation sequence compared to THOR-Net. With respect to THOR-Net, our approach reconstructs more accurate hand shapes, as it better exploits hand-object kinematic correlation and depth information. These are implicitly learned while predicting both interacting shapes simultaneously. In Table 2, we show an improvement of $0.53$ cm in hand pose estimation compared to benchmark results on the DexYCB dataset. Figure 6 also shows the qualitative reconstruction results on DexYCB without the additional refinement network.
Method Joint err. (cm) Mesh err. (cm)
A2J xiong2019a2j 2.55 -
Spurr et al. Spurr2020 2.27 -
@@
Table 2: Root-relative hand pose estimation results on $\mathcal{D}^{eval}$ of DexYCB dexycb dataset (S1) in comparison to benchmark results.
@@
We found that topological consistency is the key factor allowing $ShapeGraFormer$ to predict smooth vertices point clouds across all sequences, without the need for additional smoothness constraints, see Figure 4. After sphere-based registration, all objects share the same topology and number of vertices. We believe our choice for the sphere resolution to be a good trade-off between approximation quality and the number of vertices. Sphere-based approximation implicitly repairs irregularities in the target object shapes, making it best suited for hand-object prediction, at the cost of over-smoothed sharp edges and tiny surface details, see Figure 2.
We evaluate object reconstruction on $\mathcal{D}^{valid}$ and $\mathcal{D}^{eval}$ of HO-3D dataset, see Figure 7 and Table 3, as well as on the DexYCB dataset, see Figure 6. The additional hand-object refinement step improved object reconstruction error on all objects. This suggests that the refinement network utilizes hand information in order to improve object reconstruction. We notice that even in presence of inaccurate pose prediction as input, our approach recovers smooth object shapes, see Figure 4. Compared to THOR-Net our approach tend to oversmooth object edges, see Figure 5, possibly due to the simplified reconstruction approach.
Figure 7: Object reconstruction error (in mm) on different objects from $\mathcal{D}^{valid}$ and $\mathcal{D}^{eval}$ of HO-3D. The results show that adding a refinement GraFormer improves object reconstruction.
@@
MPVPE (cm) 5.9 6.4 7.0 9.7 11.5
Table 3: Object Reconstruction MPVPE on a selected set of objects from HO-3D and DexYCB.
4.3 Ablation Study
To validate our design choices for the $ShapeGraFormer$, we conduct an ablation study to study the modality of $\mathcal{F}_{1715}^{H}$ and the choice of layers within the GraFormer and study the impact on hand reconstruction. The ablation study in Table 4 shows that combining voxelized depth $V_{D}$, $\mathcal{F}_{V}$ and $\hat{V}^{H}$ as input to the $ShapeGraFormer$ yields the best results. Furthermore, the mixture of graph layers with transformers in the design of the GraFormer is critical to achieve best performance.
Experiment Joint err. (cm) Mesh err. (cm)
@@ 4.3 Ablation Study
Table 4: Top: An ablation study on the modality of $\mathcal{F}_{1715}^{H}$. Bottom: An ablation study on the GraFormer design choice.
@@
We test the impact of differentiable contact loss, consisting of an attraction $\mathcal{L}_{attraction}$ and a repulsion $\mathcal{L}_{repulsion}$ term, similar to Hasson et al. hasson2019learning . $\mathcal{L}_{attraction}$ is aimed at enforcing hand-object contact, by penalizing the distance between the object and the fingertips, while $\mathcal{L}_{repulsion}$ penalizes mesh interpenetration. As shown in Table 1, the additional loss term does not introduce a tangible increase in performance. Thus, suggesting our independent Refinement GraFormer component implicitly and successfully learns valid hand-object interactions.
5 Conclusion
In this paper, we propose one of the first methods for realistic hand-object pose and shape reconstruction from a single depth map. We introduce a novel 3D voxel-based GraFormer network pipeline, which reconstructs detailed 3D shapes via direct regression of mesh vertices. We conduct an ablation study to show the effectiveness of our design choices and the impact of utilizing the power of GCNs along with Transformers for hand-object shape estimation and refinement. We perform quantitative and qualitative analysis on the HO-3D dataset hampali2020honnotate and DexYCB dataset dexycb and show outstanding comparative results with the state-of-the-art. In future work, we plan to address limitations such as inaccurate annotations. In addition, we will study RGB+D methods to utilize the extra features found in RGB frames.
Acknowledgments
This work was partially funded by the Federal Ministry of Education and Research of the Federal Republic of Germany (BMBF), under grant agreements: DECODE [Grant Nr 01IW21001], GreifbAR [Grant Nr 16SV8732], and by the EU project FLUENTLY [Grant Nr 101058680].

### v2

@@ III-C VoxelNet: 3D Voxelized Shape
III-D ShapeGraFormer
To obtain topologically-coherent hand-object mesh representation, from the voxelized hand and object predictions obtained from $VoxelNet$, we train a new $ShapeGraFormer$ [41]. The $ShapeGraFormer$ includes Graph Convolutional layers and Multi-headed Attention layers to convert depth features into a pose and a shape for both the hand and the object. Although it was originally designed to lift 2D poses to 3D, GraFormer’s design combines the advantages of Graph Convolutional Networks and Transformers making it capable of effectively solving any problem that can be represented as a graph. Both the hand and the spherical objects have consistent topology and hence can be represented as a graph. Therefore, in our method, we utilize three separate GraFormers, namely: hand GraFormer, object GraFormer, and refinement GraFormer, with two feature extractors: a hand feature extractor and an object feature extractor as described in the following subsection.
@@ III-D ShapeGraFormer
III-D1 Feature Extractor and Graph Initialization
For the graph initialization, we use the outputs of the $VoxelNet$ and the raw voxelized depth as shown in Fig. 1. Namely, we extract feature maps $\mathcal{F}_{V}$ from the $VoxelNet$ and pass them through an MLP that converts them into a feature vector of size $256$. Additionally, we utilize a simple 3D CNN to extract features from the voxelized shapes and reduce them to $128$ features. Furthermore, a 3D Max Pool layer reduces the size of the raw voxelized depth from $44\times 44\times 4$ 44 to $11\times 11\times 1$ 11 creating a $1331$ feature vector. These features are then combined to create a hand feature vector $\mathcal{F}^{H}_{1715}$ of size $1715$ to initialize the graph vertices as shown in Fig. 1. The same operation is repeated for the object’s vertices resulting in $\mathcal{F}^{O}_{1715}$. To initialize the adjacency matrix of the graph layers, we use the mesh faces of the MANO model along with the faces of the spherical mesh as described in Section III-A.
III-D2 Positional Embedding
In order to generate distinct features for each vertex in the graph, we propose a positional embedding layer that converts the vertices of the template meshes for both the hand and object into positional vectors $\mathcal{E}_{i}^{p}$ of the same size as the feature vector where i is the index of the ith vertex in the combined hand-object graph. The positional vectors $\mathcal{E}_{i}^{p}$ are then accumulated with the shared feature vectors $\mathcal{F}^{H}_{1715}$ and $\mathcal{F}^{O}_{1715}$ depending on whether the ith vertex belongs to the hand or the object in order to create a unique representation for each vertex. For the hand mesh, we adopt the default MANO hand as the template, while for the object mesh, we use the sphere that is utilized for deforming objects, see Section III-A.
III-D3 GraFormer Details
Each GraFormer consists of five consecutive GraAttention components followed by Chebyshev Graph Convolutional layers as described in [41] and shown in Fig. 3. A GraAttention component consists of a Multi-headed Attention layer where the hidden dimension is $128$ and the number of heads is $4$, following the ablation study mentioned in [41] where they studied the impact of hidden dimension size and number of layers in the GraFormer on pose estimation. Compared to normal Transformers, the last fully connected layer of the Multi-headed Attention in the GraFormer is a graph convolutional layer, not a feed-forward layer. The GraFormer also contains an input layer that maps the feature vector to the hidden dimension size and an output layer that maps the hidden dimension into the corresponding 3D coordinate value for each vertex in the graph.
@@ III-D3 GraFormer Details
III-D4 Refinement GraFormer
To enhance the realistic appearance and correct minor shape-related artifacts in the predicted object shapes, we add another GraFormer for refinement. We utilize the initial shape produced by the Hand and Object GraFormer as input to another positional embedding layer while we use $\mathcal{F}^{H}_{1715}$ and $\mathcal{F}^{O}_{1715}$ to initialize the new graph. In the case of using a refinement GraFormer during training, we get $2$ separate shapes from the $ShapeGraFormer$ and apply the same loss function mentioned in III-D on both of them.
III-E Training and Implementation Details
We train all the components of our network on fully annotated public hand-object pose and shape datasets, namely, HO-3D [28] and DexYCB [29]. For training, we use the Adam [43] optimizer with a learning rate set to .001.001.001.001. For improved convergence, we train $PoseNet$ separately, then fix the obtained weights to train the remaining network components. All learning and inference are implemented in PyTorch and are conducted on an NVIDIA A100 GPU.
IV Experiments
In this section, we quantitatively and qualitatively evaluate the effectiveness of our 3D hand-object reconstruction pipeline on two popular datasets, namely HO-3D dataset (version v2 and v3) [28] and DexYCB [29]. We additionally compare our pipeline with state-of-the-art approaches in Section IV-B. For quantitative evaluation and comparisons, we use the following two metrics: (i) the average 3D joint location error ($\mathcal{J}$ err.), and (ii) the average 3D vertex location error ($\mathcal{V}$ err.) over all test frames.
IV-A Datasets
@@
The HO-3D dataset [28], is a publicly available dataset with 3D pose annotations for hands interacting with objects captured from third-person views. The dataset has multiple versions and we report the results on v2 and v3 of the dataset. For HO-3D (v3), the training set $\mathcal{D}:=\{\mathcal{D}^{train},\mathcal{D}^{valid}\}$ contains annotations for $|\mathcal{D}^{train}|=71,662$ and $|\mathcal{D}^{valid}|=10,927$ images and a total of $55$ sequences and $10$ different objects from the YCB-Video dataset [44], $9$ for the training set and $1$ (unseen) for the evaluation set. The evaluation set $\mathcal{D}^{eval}$ comprises $13$ sequences with a total of $|\mathcal{D}^{eval}|=20,137$ frames addressing challenging scenarios: namely, (i) $3$ sequences with ($2$) seen objects and seen hands, (ii) $5$ sequences with ($1$) seen object but unseen hands, and (iii) $5$ sequences with seen hands but $1$ unseen object. Hands in the evaluation set are just annotated with the wrist coordinates, while the full hand is not annotated. Object pose and shape are annotated over all available sets.
@@
To extend our evaluation, we also train our network on the DexYCB dataset [29]. DexYCB contains hand pose and shape and 6D object pose annotations for  $582k$ frames recorded on $10$ different subjects, using $20$ different objects and $8$ views. We use the S1 evaluation setup as specified by the authors where $\mathcal{D}^{valid}$ contains $1$ unseen subject and $\mathcal{D}^{eval}$ contains $2$ unseen subjects and $\mathcal{D}^{train}$ contains 7 subjects and all different objects. The exact split sizes are: $|\mathcal{D}^{train}|=407,088$, $|\mathcal{D}^{valid}|=58,592$, and $|\mathcal{D}^{eval}|=116,288$.
@@
IV-B Evaluation
In this section, we evaluate the hand-object shape and pose reconstruction and compare it to state-of-the-art approaches in challenging scenarios.
@@ IV-B Evaluation
We compare our work on HO-3D quantitatively with six different methods. The work of Hasson et al. [23] and THOR-Net  [20] are the most related to ours in terms of the goals. However, as they focus on RGB inputs, comparisons are made up to scale. Malik et al. [9] on the other hand, is based on depth inputs and has a comparable voxel-based network pipeline to our design. However, they ignore the presence of an object and reconstruct hands in isolation. We include the representative RGB-based hand reconstruction approach of Hampali et al. [28], and ArtiBoost  [31] for completeness. For a fair comparison, all methods have been (re-)trained on the HO-3D dataset. The first two methods provide publicly available results (hand-only), which we report in Table I. We re-implemented HandVoxNet following the authors’ instructions and trained all the network components on HO-3D. In addition to that, we also report the root-relative pose estimation error in Table II and compare it to two benchmark methods mentioned by the DexYCB authors. We also show qualitative samples for hand reconstruction from DexYCB in Fig. 6.
@@
Quantitative evaluation on hand pose and shape reconstruction, shows that our method outperforms existing methods, see Table I and Figures 4, 5.
Method $\mathcal{J}$ err. (cm) $\mathcal{V}$ err. (cm)
Hampali et al. [28] 8.42 8.34
Hasson et al. [23] 4.97 4.96
HandVoxNet [9] 2.77 2.67
HandVoxNet++ [40] 2.46 2.70
THOR-Net [20] 2.63 2.63
@@
ArtiBoost [31] 2.26 10.95
THOR-Net [20] 2.56 2.37
@@
TABLE I: Quantitative results on the evaluation set $\mathcal{D}^{eval}$ of HO-3D [28] dataset. Upper table: HO-3D (v2), lower table: HO-3D (v3).
We achieve a minimum pose and shape prediction improvement respectively of $0.27$ cm and $0.43$ cm over the state-of-the-art. The experiments show that training the network for hand reconstruction alone without object reconstruction improves hand reconstruction and that the refinement stage has no impact on hand reconstruction as shown in Table I. However, joint training for the hand and the object outperforms other methods. Fig. 5 shows our results on a few frames of the challenging evaluation sequence compared to THOR-Net. In comparison to THOR-Net, our approach reconstructs more accurate hand shapes, as it better exploits hand-object kinematic correlation and depth information. These are implicitly learned while predicting both interacting shapes simultaneously. In Table II, we show an improvement of $0.53$ cm in hand pose estimation compared to benchmark results on the DexYCB dataset. Fig. 6 also shows the qualitative reconstruction results on DexYCB without the additional refinement network.
Method $\mathcal{J}$ err. (cm) $\mathcal{V}$ err. (cm)
A2J [45] 2.55 -
Spurr et al. [46] 2.27 -
@@
TABLE II: Root-relative hand pose estimation results on $\mathcal{D}^{eval}$ of DexYCB [29] dataset (S1) in comparison to benchmark results.
@@
We found that topological consistency is the key factor allowing $ShapeGraFormer$ to predict smooth vertices point clouds across all sequences, without the need for additional smoothness constraints, see Fig. 4. After sphere-based registration, all objects share the same topology and number of vertices. We believe our choice for the sphere resolution to be a good trade-off between approximation quality and the number of vertices. Sphere-based approximation implicitly repairs irregularities in the target object shapes, making it best suited for hand-object prediction, at the cost of over-smoothed sharp edges and tiny surface details, see Fig. 2.
We evaluate object reconstruction on $\mathcal{D}^{valid}$ and $\mathcal{D}^{eval}$ of HO-3D dataset, see Fig. 7 and Table III, as well as on the DexYCB dataset, see Fig. 6. The additional hand-object refinement step improved object reconstruction error on all objects. This suggests that the refinement network utilizes hand information to improve object reconstruction. We notice that even in the presence of inaccurate pose prediction as input, our approach recovers smooth object shapes, see Fig. 4. Compared to THOR-Net our approach tends to oversmooth object edges, see Fig. 5, possibly due to the simplified reconstruction approach.
Figure 7: Object reconstruction error ($\mathcal{V}$ err. in mm) on different objects from $\mathcal{D}^{valid}$ and $\mathcal{D}^{eval}$ of HO-3D. The results show that adding a refinement GraFormer improves object reconstruction.
@@
$\mathcal{V}$ err. (cm) 5.9 6.4 7.0 9.7 11.5
TABLE III: Object’s reconstruction error ($\mathcal{V}$ err. (cm)) on a selected set of objects from HO-3D and DexYCB.
IV-C Ablation Study
To validate our design choices for the $ShapeGraFormer$, we conduct an ablation study to study the modality of $\mathcal{F}_{1715}^{H}$ and the choice of layers within the GraFormer and study the impact on hand reconstruction. The ablation study in Table IV shows that combining voxelized depth $V_{D}$, $\mathcal{F}_{V}$, and $\hat{V}^{H}$ as input to the $ShapeGraFormer$ yields the best results. Furthermore, the mixture of graph layers with transformers in the design of the GraFormer is critical to achieving the best performance.
Experiment $\mathcal{J}$ err. (cm) $\mathcal{V}$ err. (cm)
@@ IV-C Ablation Study
TABLE IV: Top: An ablation study on the modality of $\mathcal{F}_{1715}^{H}$. Bottom: An ablation study on the GraFormer design choice.
@@
We test the impact of differentiable contact loss, consisting of an attraction $\mathcal{L}_{attraction}$ and a repulsion $\mathcal{L}_{repulsion}$ term, similar to Hasson et al. [23]. $\mathcal{L}_{attraction}$ is aimed at enforcing hand-object contact, by penalizing the distance between the object and the fingertips, while $\mathcal{L}_{repulsion}$ penalizes mesh interpenetration. As shown in Table I, the additional loss term does not introduce a tangible increase in performance. Thus, suggesting our independent Refinement GraFormer component implicitly and successfully learns valid hand-object interactions.
V Conclusion
In this paper, we propose one of the first methods for realistic hand-object pose and shape reconstruction from a single depth map. We introduce a novel 3D voxel-based GraFormer network pipeline, which reconstructs detailed 3D shapes via direct regression of mesh vertices. We conduct an ablation study to show the effectiveness of our design choices and the impact of utilizing the power of GCNs along with Transformers for hand-object shape estimation and refinement. We perform quantitative and qualitative analysis on the HO-3D dataset [28] and DexYCB dataset [29] and show outstanding comparative results with the state-of-the-art. In future work, we plan to address limitations such as inaccurate annotations. In addition, we will study RGB+D methods to utilize the extra features found in RGB frames.